OpenStudy (anonymous):

Algebra 1 - Please explain how to get the correct answer? Medal goes to best explained answer!! Michelle rents a movie for a flat fee of $1.50 plus an additional $1.25 for each night she keeps the movie. Choose the cost function that represents this scenario if x equals the number of nights Michelle has the movie. c(x) = 1.50 + 1.25x c(x) = 1.50x + 1.25 c(x) = 2.75 c(x) = (1.50 + 1.25)x

OpenStudy (anonymous):

think about which one is applied once and which one needs to go up for every night

OpenStudy (anonymous):

@lonnie455rich Okay so the amount that will remain the same is $1.50, and the increasing/changing amount is $1.25 Where does that leave me in this problem.. because my guestimate is that the answer is the last option.

OpenStudy (amigatour):

C(x)=1.50+1.25x We know that the flat fee is $1.50, which is the cost of the movie. The $1.25 is a charge for an unknown nights she keep the movies. If she doesn't return it on the day she has to, she will be charged of $1.25/night. Example: She didn't have an over due. C(0)=1.50+1.25(0) =1.50

OpenStudy (amigatour):

So, if she kept her movie for a week, it will be like this C(7)= 1.50+1.25(7).

OpenStudy (the_fizicx99):

Its a flat fee of $1.50 + 1.25x for every movie she rents You can make a function: C(x) = 1.50 + 1.25x C(1) = 1.50 + 1.25*1 = 2.75 Where x is the number of movie she rents.

OpenStudy (the_fizicx99):

You can plug in anything for x. C(50) = 1.50 + 1.25*50 1.50 + 62.5 = $64 :)

OpenStudy (amigatour):

It's hard for me to explain it in English. Lol. I will try. However, pretend you rent movies on December 06, 2013. The lady told you that the Dark Night disc cost $1.50 to rent. You have to return your movie on Dec. 08, 2013. If you don't return it before or on 12/8/13, you will be charged $1.25 per night. Here, we don't know how many nights you have kept the movie, so we label it as X. That's why the function is C(x)=1.50+ 1.25x

OpenStudy (amigatour):

I hope that help.

OpenStudy (anonymous):

@amigatour so it is the first option?

OpenStudy (the_fizicx99):

Yes.

OpenStudy (amigatour):

yes, it is.

OpenStudy (amigatour):

You can do it at f(x)=1.25x+1.50 It doesn't matter as long you get the correct function. Some professors/teachers are picky, so follow their instruction. They want you to write it like this f(x)=1.25x+1.50

OpenStudy (anonymous):

@amigatour okay thanks so much! :)